Out of 5 Mathematicians and 7 physicists, a committee consisting of 2 mathematicians and 3 physicists is to be formed. In how many ways can this be done if.
a) any mathematician and any physicists can be included
b) one particular physicist must be on the committee
c) 2 particular mathematicians cannot be on the committee

Answer by Boreal(15235)

a: 5C2*7C3=10*35=350
b. 5C2*6C2*1=150, because the specific physicist takes 1 from each part.
c. 3C2*7C3=3*35=105